For a mole of ideal gas at T = 35 ºC, what is the work done for an isothermal expansion from a

volume V0 to 10V0


1
Expert's answer
2020-06-18T10:58:10-0400

Isothermal expansion work equals

"A=\nu\cdot R \cdot T \cdot \ln{\frac{V_2}{V_1}}"

Where

"\nu=1 mole"

"R=8.814J/mole \cdot K"

"T=273+35=308K"

"V_1=V_0"

"V_2=10V_0"

Then

"A=1\cdot 8.314 \cdot 308 \cdot \ln{\frac{10V_0}{V_0}}=2561 \cdot\ln{(10)}=5896J"